TBASERDN040213 February 13, 2004 Q-band receiver in the Equipment Room, cold. Using lab 83620 as LO1, set at 9250MHz. Connected IF of L1 and L2 to OR 1 and 3 respectively, by coax. These tests are after removing LO multipliers found with bias oscillations L1 -> OR1 -> CM1 -> SF1 -> ACS J1 (bank 1, rcvr 1) cM2 -> SF2 -> ACS J3 (bank 2, rcvr 1) L2 -> OR3 -> CM5 -> SF5 -> ACS J2 (bank 1, rcvr 2) CM6 -> SF6 -> ACS J4 (bank 2, rcvr 2) T=120s, integration=10s. ACS in two bank, two sampler/bank mode. Scans 1-6 w/ room temp. load Scans 7-16 w/ LN2 load. Changed R1->OR1; R2->OR3. Scans 17-25 LN2 load Adjusted bias on R2: VD1->1.50 from 0.882 (VG1-> +0.31) Scans 26-34 Adjusted bias on R2: VD1->1.70 from 1.50 (VG1->-0.01 Scans 35-42 Changed to L1 and L2, rebalanced spectrometer. Scans 49-50 (system crashed during 50). -Break- Resumed ~`20:00:00 UT Adjusted first stage all four chs to 1.70V, 1.0mA. L1->OR1, L2->OR3 Scans 101-108. NT high. Adjusted bias L1 and L2 -> 1.2V, 1.5mA Scans 110-114 Changed to R1->OR1, R2->OR3. (1.70V, 1.0mA) Scans 115-119. Changed bias R1,R2 to 1.2V,1.5mA. Scans 120-129 Analysis: See the Tex report in this directory.